OpenStudy (anonymous):

Can someone please help me!? ASAP!!!!! A 1-pound weight has a mass of about 0.45 kilogram. What is the mass in kilograms of a sculpture that weights 570 pounds? Round your answer to the nearest tenth of a kilogram.

OpenStudy (anonymous):

256.6

OpenStudy (anonymous):

or, 257 pounds

OpenStudy (anonymous):

How did you get it? @Bandit56

OpenStudy (anonymous):

times 0.45 by 570

OpenStudy (anonymous):

ohhh it kilograms not pounds, so 257 kilograms

OpenStudy (anonymous):

Wouldn't the answer be 256.6kg? @Bandit56

OpenStudy (jdoe0001):

\(\large \begin{array}{llll} lbs&kgs\\ \hline\\ 1&0.45\\ x&570 \end{array}\implies \cfrac{1}{x}=\cfrac{0.45}{570}\)

OpenStudy (anonymous):

Is my answer correct @jdoe0001

OpenStudy (jdoe0001):

ohhh.... it was 570 lbs.. so... ok then \(\large \begin{array}{llll} lbs&kgs\\ \hline\\ 1&0.45\\ 570&x \end{array}\implies \cfrac{1}{570}=\cfrac{0.45}{x}\)

OpenStudy (jdoe0001):

Wouldn't the answer be 256.6kg \(\large \checkmark\)

OpenStudy (jdoe0001):

\(\large \begin{array}{ccllll} lbs&kgs\\ \hline\\ 1&0.45\\ 570&x \end{array}\implies \cfrac{1}{570}=\cfrac{0.45}{x}\implies x=570\times 0.45\)